OpenStudy (anonymous):

Use the binomial coefficient formula to calculate the binomial coefficient for the 3rd term in a polynomial expansion when the power of the binomial is 5.

OpenStudy (anonymous):

good quetion

OpenStudy (anonymous):

i could not get right answer..it is multiple choice

OpenStudy (anonymous):

could u check answer ..... Use the binomial coefficient formula to calculate the binomial coefficient for the 4th term in a polynomial expansion when the power of the binomial is 6. answer 20

OpenStudy (anonymous):

\[\left(\begin{matrix}6 \\ 3\end{matrix}\right)x^3y^3= 20x^3y^3\]

OpenStudy (anonymous):

The 3rd term of power 5 is: \[\left(\begin{matrix}5 \\ 2\end{matrix}\right)x^3y^2= 10 x^3y^2\]